OpenStudy (anonymous):

solve 3(4x − 2x) = 8

OpenStudy (darkigloo):

first you have to distribute the 3.

OpenStudy (darkigloo):

Then combine like terms on the left side.

OpenStudy (darkigloo):

did you do that?

OpenStudy (anonymous):

yes i did and i didnt get the answer that were already being shown

OpenStudy (darkigloo):

\[3(4x-2x)=8\] 12x - 6x = 8 6x = 8 x = 8/6